OpenStudy (darkigloo):

If y= e^(3ln(x^2)) find y'

OpenStudy (darkigloo):

the answer is 6x^5 but i don't know how.

OpenStudy (dumbcow):

simplify first \[e^{3 \ln(x^2)} = (e^{\ln(x^2)})^3 = (x^2)^3 = x^6\]

OpenStudy (darkigloo):

ohh. so then you find the derivative of x^6 which is 6x^5. thank you.

OpenStudy (dumbcow):

no problem